# Latent Topic Synthesis: Leveraging LLMs for Electoral Ad Analysis

### v1

@@ Introduction
To identify the moral foundation of each clusters, we first prompt the LLM to summarize the primary talking point of the cluster representatives with regards to the annotated label. The LLM is then provided with definitions of the six moral foundations, and the extracted argument is classified using constrained decoding to determine the most strongly aligned foundation. Prompts are provided in the Appendix.
Supervised Classification
To annotate the unlabeled ads, we use the labeled cluster representatives as the training set for a supervised classification task. We used a variety of classifiers, including Logistic Regression, XGBoost (Chen and Guestrin 2016), RoBERTa (Liu et al. 2019), and SetFit (Tunstall et al. 2022).
@@
For each model, we used the all-MiniLM-L6-v2 model to embed the ad texts and employed GridSearchCV to optimize the hyperparameters. We used $5$-fold cross-validation to evaluate the models, and selected the best performing model based on the macro F1 score. The results of this evaluation are shown in Table 3.
@@
Demographic Targeting Analysis
Using the ad impressions data, we can analyze the targeting data of the ads. The analysis focuses on the use of social media platforms, particularly Meta and Instagram, to disseminate targeted political advertisements. We examine the strategies employed by the advertisers to reach specific demographic groups.
Meta allows advertisers to define their audiences on the basis of location (state-level), age and gender. We utilized this data, together with the annotations, to analyze the microtargeting strategies of the political entities.
To understand how messages are tailored to different demographics, we looked at the positive pointwise mutual information (PPMI) between the words in the advertisements and the demographic groups. The PPMI is a measure of association between two events, in this case, the words in the advertisements and the demographic groups. A higher PPMI value indicates a stronger association between a word and a demographic group. We compared both the topics discussed in the advertisements as well as the moral foundations that were used to appeal to the different demographic groups. This annotated corpus enabled detailed analysis of issue emphasis and demographic targeting, which we report in the Results and Analysis Section.
Results and Analysis
We evaluate the performance of our approach on the social media advertisement dataset, evaluating the performance of both annotation and classification tasks. Additionally, we discuss the results of our case study, showcasing differences in moral framing and issue distribution across different demographics.
@@
Figure 2: Average spend and impressions reveal which topics dominate the ad landscape.
Figure 3: Top five funding entities by total spend across economy, abortion, and crime/justice topics. Economic ads are dominated by conservative groups, abortion spending reflects polarized investments from both anti-abortion and pro-abortion rights organizations, and crime/justice advertising is driven by a mix of PACs and local campaigns.
Baseline Comparison
We evaluated the topic labels, comparing the performance of our proposed framework with the baseline BERTopic  (Grootendorst 2022). For both models, the same clusters were used, and both models used the same LLM to generate the topic labels. For BERTopic, the LLM generated the labels based on both the representative documents and the top-10 keywords.
Two annotators were each given 10 randomly selected clustered ads. The annotators were volunteer graduate students who spoke fluent English with backgrounds in Computational Social Science and Computer Science. For each ad, they were given both the label from BERTopic and our proposed framework (the order of the labels was randomized). The annotators were asked to select the label that best represented the ad, and to provide a score from 1-5 for each label, where 5 would be a perfect fit, and 1 would describe a totally irrelevant label. The results of this comparison are shown in Table 4.
These results show that our proposed framework outperforms BERTopic in terms of both average score and the number of times the label was selected as the best label. A Cohen’s Kappa (Cohen 1960) score of $0.66$ was achieved between the two annotators, indicating a moderate level of agreement. For $5$ out of the $20$ rounds, the annotators gave the same score for both models (in all cases a score of 1). The average score of $2.8$ for our proposed framework indicates that the labels generated are more relevant and accurate compared to the average score of $1.2$ for BERTopic.
Spending and Reach by Topic
We begin with a descriptive analysis of ad spending and exposure across topics. Figure 2 shows the average spend (left panel) and average impressions (right panel) per topic. Overall, voting and immigration dominate both spending and impressions, while issues such as healthcare access and property taxes receive relatively little paid attention. Interestingly, abortion and election integrity attract intermediate spending yet generate disproportionally high impressions, suggesting higher engagement or lower cost per impression in these issue domains.
Figure 4: Topic-specific moral narratives: multiple frames shape economic ads, Liberty/Oppression dominates abortion related ads.
Top Funders by Issue Domain
We examine the concentration of spending among funding entities. Figure 3 presents the top $5$ funding entities by total spend for one of the three central issues [^2] https://www.pewresearch.org/politics/2024/09/09/issues-and-the-2024-election/ in the 2024 election: economy, abortion, and crime/justice.
On the economy, conservative-aligned groups dominate the funding landscape. The Keystone [^3] https://www.keystonepac.org/ and the Congressional Leadership Fund [^4] https://congressionalleadershipfund.org/ lead spending, followed by organizations such as Americans for Prosperity Action [^5] https://americansforprosperity.org/. This pattern reflects Republican priorities around fiscal and economic messaging, emphasizing taxation, regulation, and inflation.
For abortion, spending is highly polarized. Florida Voters Against Extremism [^6] https://dos.elections.myflorida.com/committees/ComDetail.asp?account=84315 is the largest spender, anchoring anti-abortion rights mobilization, particularly around state-level ballot initiatives. In contrast, progressive-aligned group such as Montanans Securing Reproductive Rights  [^7] https://www.influencewatch.org/organization/montanans-securing-reproductive-rights/ invests heavily in support of abortion rights. This split underscores abortion as a wedge issue, with both sides committing substantial but asymmetrically distributed resources.
In the case of crime and justice, the funding environment is distinct. Judicial Fairness PAC [^8] https://judicialfairnesspac.org/ (with messages focused on judicial appointments and criminal justice reform) and Smart and Safe Florida [^9] https://smartandsafeflorida.com/ (with messages focused on public safety) together account for a large share of spending. Other contributors, such as Stop Houston Murders PAC [^10] https://stophoustonmurders.com/ and Iowans for Zach Nunn [^11] https://www.fec.gov/data/committee/C00784389/, highlight localized campaign investments. Notably, this issue domain demonstrates geographically fragmented but financially intensive spending, reflecting its resonance in state and local politics as well as in national-level partisan framing.
Overall, Figure 3 shows that while the economy and abortion are dominated by major national organizations with strong partisan alignment, crime/justice advertising is more decentralized, fueled by a mix of PACs and candidate-aligned committees targeting specific jurisdictions.
Moral Foundations in Funded Messaging
We link moral framing to financial investment. Figure 4 compares the top $5$ funding entities in two salient topics — economy and abortion — broken down by their dominant moral foundation appeals. On the economy, funding entities split their resources between Care/Harm frames (e.g., worker hardship) and Fairness/Cheating or Liberty/Oppression frames (e.g., taxation, regulation). On abortion, almost all spending is channeled through a Liberty/Oppression frame, consistent with rhetoric around government interference in reproductive rights. This illustrates that moral narratives are not evenly distributed across topics: economic appeals are morally plural, while abortion appeals are morally singular.
Topic Salience and Moral Framing
We present some selected results from our case study. Firstly, we analyze the correlation between moral foundations and the latent topics, taking LDA topics as a baseline. To compute the correlation, we first one-hot encoded each ad twice: once by topic and once by moral foundation. We then computed the pairwise correlation using the Pearson correlation between these one-hot encodings. The results are shown in Figure 5. In Figure 5(a), we can see very strong correlations between certain moral foundations and topics. For example, the Fairness/Cheating moral axis is strongly correlated with the crime/justice topic, while the Loyalty/Betrayal moral axis is strongly correlated with the immigration topic. This suggests that our proposed framework is able to capture the moral framing of the advertisements. In contrast, the LDA topics in Figure 5(b) show much weaker correlations with the moral foundations.
@@
(a) Topics: Females in Alabama vs Washington
(b) Topic: All Age groups in Florida
(c) Topic: Males in Montana vs Virginia
(d) Moral Framing across States
We also analyzed the distribution of moral framings and political issues across different demographics to uncover political microtargeting strategies. We used the positive point-wise mutual information (PPMI) to detect topics that appear disproportionately in ads shown to specific demographics, potentially indicating targeted messaging. We visualize the results of the PPMI analysis in Figure 6. The heatmaps show the distribution of moral framings and political issues across different demographics. Some interesting observations include the differences in ads targeting different states, or the shift in topics among different age groups in Florida (Figure 6(b)). For example, young people in Florida are more likely to be targeted with ads related to affordable housing, while older people are more likely to be targeted with ads related to abortion. This shows generational targeting strategies, where economic concerns are emphasized for younger groups, while older groups are mobilized on moralized issues like abortion.
Conclusion
We produce a modular framework for large-scale annotation of unstructured text documents. While our framework is generalizable, the key contribution here is enabling large-scale, interpretable analysis of electoral ads. Applied to a corpus of eight thousand political advertisements, our method enabled efficient labeling with minimal human effort and low computational cost. We demonstrate that our approach is effective for both supervised and unsupervised learning tasks, and the analysis of the resulting annotations uncovered meaningful patterns in issue framing across demographic groups. In addition, the release of a new dataset will provide a valuable resource for the research community. We believe our framework can be applied to a variety of text classification tasks, discovering latent structures in large corpora, and enable cheap and efficient automated annotation of unstructured documents. Our findings demonstrate that combining clustering with LLM-based labeling yields scalable, interpretable analysis of political ads, surfacing both substantive issue agendas and the moral frames used to mobilize audiences. These insights show how social media advertising both reflects and amplifies polarization, providing researchers and policymakers with tools to monitor emerging narratives and targeting strategies.
Limitations
@@ Conclusion
While our framework reduces manual annotation effort and generalizes across domains, several limitations remain. Firstly, we purposefully choose a weaker LLM to show that our framework is not limited to computational resources or budget constraints. However, the performance of our framework is likely to improve with more powerful LLMs. Secondly, while we demonstrate the effectiveness of our framework on a political advertisement corpus, it may not generalize to all domains.
@@ Appendix A Prompts
Can the ads can be summarized by one of the previous topics: Answer with ”yes” or ”no”.
@@
In three words or less, describe the issue that ads discuss, summarizing the topic. Examples: ”Abortion”, ”Climate Change”, etc.
Output constriained to [”yes”, ”no”].
@@
A better topic for these ads is: ”
The LLM would finish the above, with the quote (”) as a stop token.
Annotation

### v2

@@ 4.3 Moral Foundations
To identify the moral foundation of each cluster, we first prompt the LLM to summarize the primary talking point of the cluster representatives with regard to the annotated label. The LLM is then provided with definitions of the six moral foundations, and the extracted argument is classified using constrained decoding to determine the most strongly aligned foundation. Prompts are in App. A.
4.4 Supervised Classification
To annotate the unlabeled ads, we use the labeled cluster representatives as the training set for a supervised classification task. We used a variety of classifiers, including Logistic Regression, XGBoost Chen and Guestrin (2016), RoBERTa Liu et al. (2019), and SetFit Tunstall et al. (2022).
@@ 4.4 Supervised Classification
For each model, we used the all-MiniLM-L6-v2 model to embed the ad texts and employed GridSearchCV to optimize the hyperparameters. We used $5$-fold cross-validation to evaluate the models, and selected the best-performing model based on the macro F1 score. The results of this evaluation are shown in Table 3.
@@
4.5 Demographic Targeting Analysis
Using the ad impressions data, we can analyze the targeting data of the ads. The analysis focuses on the use of social media platforms, particularly Meta and Instagram, to disseminate targeted political advertisements. We examine the strategies employed by the advertisers to reach specific demographic groups. Meta allows advertisers to define their audiences on the basis of location (state-level), age and gender. We utilized this data, together with the annotations, to analyze the microtargeting strategies of the political entities.
To understand how messages are tailored to different demographics, we look at the positive pointwise mutual information (PPMI) Church and Hanks (1990) between the words in the advertisements and the demographic groups. The PPMI is a measure of association between two events, in this case, the words in the advertisements and the demographic groups. A higher PPMI value indicates a stronger association between a word and a demographic group. We compare both the topics discussed in the advertisements and the moral foundations that are used to appeal to the different demographic groups. This annotated corpus enables detailed analysis of issue emphasis and demographic targeting, reported in Section 5.
5 Results and Analysis
We evaluate the performance of our approach on the social media advertisement dataset, evaluating the performance of both annotation and classification tasks. Additionally, we discuss the results of our case study, showcasing differences in moral framing and issue distribution across different demographics. Our analysis proceeds at two complementary levels: (i) cluster-level analysis, which characterizes issue salience and moral framing using LLM-synthesized topic clusters, and (ii) ad-level analysis, which leverages individual ad annotations to examine spending patterns, funding style, and moral framing at scale.
@@ 5 Results and Analysis
5.1 Baseline Comparison
First, we evaluate the topic labels by comparing the performance of our proposed framework with the baseline BERTopic  Grootendorst (2022). For both models, the same clusters are used, and both models use the same LLM to generate the topic labels. For BERTopic, the LLM generated labels based on both the representative documents and the top 10 keywords.
To assess whether iterative taxonomy construction provides benefits beyond direct LLM labeling, we introduce an additional baseline inspired by TopicGPT Pham et al. (2023). For each cluster, we prompt the same LLM with the cluster representative ads and ask it to generate a concise topic label in a single step, without access to previously generated labels or iterative refinement. This baseline isolates the effect of structured taxonomy growth and constrained validation in our framework.
Two annotators were each given $10$ randomly selected clustered ads. The annotators were volunteer graduate students and faculty who spoke fluent English with backgrounds in computational social science, natural language processing, and computer science. For each ad, they were given both the label from BERTopic and our proposed framework (the order of the labels was randomized). The annotators were asked to select the label that best represented the ad, and to provide a score from 1-5 for each label, where $5$ would be a perfect fit, and $1$ would describe a totally irrelevant label. The results of this comparison are shown in Table 4.
These results suggest that the proposed framework yields more consistent and better-aligned topic labels than both BERTopic and the single-shot TopicGPT-style baseline under this evaluation setup. A Cohen’s Kappa Cohen (1960) score of $0.66$ was achieved between the two annotators, indicating a moderate level of agreement. For $5$ out of the $20$ rounds, the annotators gave the same score for both models (in all cases a score of 1).
Analyses.
The following analyses are intended to illustrate the types of downstream insights enabled by the induced taxonomy, rather than to make exhaustive claims about political advertising behavior.
5.2 Topic Salience and Moral Framing
We present some selected results from our case study. Firstly, we analyze the correlation between moral foundations and the latent topics, taking LDA topics as a baseline. To compute the correlation, we first one-hot encoded each ad twice: once by topic and once by moral foundation. We then computed the pairwise correlation using the Pearson correlation between these one-hot encodings. The results are shown in Fig. 2. In Fig. 2(a), we can see very strong correlations between certain moral foundations and topics. For example, the Fairness/Cheating moral axis is strongly correlated with the crime/justice topic, while the Loyalty/Betrayal moral axis is strongly correlated with the immigration topic. This suggests that the induced topics align with moral framing patterns in this dataset. In contrast, the LDA topics in Fig. 2(b) show much weaker correlations with the moral foundations.
We analyze demographic targeting using positive PPMI between ads and audience attributes (Fig. 4 in App. B). Clear age- and location-specific patterns emerge: in Florida, younger audiences are disproportionately shown affordable housing ads, while older audiences receive more abortion-focused messaging. Geographic contrasts are also evident: male audiences in Montana are more exposed to environmental protection and personal freedom ads, whereas male audiences in Virginia are targeted more heavily with crime/justice and voting-related content, reflecting systematic alignment between issue framing and demographic context.
Figure 3: Average spend and impressions reveal which topics dominate the political ad landscape.
5.3 Ad-Level Analysis
To enable finer-grained analysis beyond cluster-level structure, we annotate individual ad using the synthesized topic taxonomy. Using the generated topics as labels, we apply an LLM-based annotation (model: mistralai/Mistral-7B-v0.1 Jiang et al. (2023)) to assign each ad a topic and moral foundation. This enables ad-level analysis of spending, exposure, and funder behavior, which we report below.
5.3.1 Spending and Reach by Topic
We begin with a descriptive analysis of ad spending and exposure across topics. Fig. 3 shows the average spend (left panel) and average impressions (right panel) per topic. Topics such as abortion, personal freedom, and voting rights receive the highest average spend, indicating strong advertiser investment in normatively salient and mobilizing issues. In contrast, immigration and border security achieve comparatively high impressions with lower average spend, suggesting broader organic reach or more cost-efficient targeting. Issues such as property taxes and local services (e.g., sewer sale) receive minimal investment and exposure. Overall, the divergence between spend and impressions highlights strategic differences in how advertisers allocate resources versus how effectively topics scale to large audiences.
5.3.2 Top Funders by Issue Domain
We analyze the concentration of advertising spend across funding entities for three central issues of 2024 US presidential election—economy, abortion, and crime/justice [^2] https://www.pewresearch.org/politics/2024/09/09/issues-and-the-2024-election/ (Fig. 5, App. C.1).
Spending on the economy is highly concentrated, with one dominant national organization and a sharp drop-off among other major PACs. Abortion exhibits the strongest concentration overall, with a small number of opposing organizations accounting for the majority of spending, underscoring its role as a high-salience wedge issue. In contrast, crime/justice shows a more distributed funding profile, with multiple PACs contributing comparable levels of investment, often tied to specific states or jurisdictions. App. C.1 provides the details. Overall, these patterns reveal substantial variation in funding concentration across issues: economy and abortion messaging are driven by a small set of major actors, whereas crime/justice advertising reflects more decentralized and localized campaign strategies.
5.3.3 Moral Framing and Ad Spending
We examine how moral framing aligns with financial investment across issue domains. Aggregate spending by moral foundation (Fig. 6, App. C.2) reveals distinct moral structures across topics. Economy advertising distributes spending across multiple foundations—primarily Fairness/Cheating and Liberty/Oppression—indicating moral plurality. In contrast, abortion advertising concentrates spending on two sharply opposed foundations, Liberty/Oppression and Sanctity/Degradation, reflecting moral polarization.
Funder-level breakdowns (Fig. 7, App. C.2) show that these aggregate patterns arise from strategic specialization by major funding entities rather than uniform framing across actors. Detail in App. C.2
6 Conclusion
We propose a general, seed-free framework for iterative topic taxonomy induction using LLMs. By combining embedding-based clustering with constrained, multi-pass LLM inference, the framework enforces global consistency across clusters and improves topic label quality relative to both classical topic models and single-shot LLM labeling under human evaluation. We validate the framework through a large-scale political advertising case study, demonstrating how the induced taxonomies enable interpretable downstream analyses such as issue prevalence, moral framing, and demographic targeting in this setting. While our empirical evaluation focuses on electoral advertising, the framework itself is not domain-specific and can be applied to other large, unlabeled text corpora. An important direction for future work is to evaluate taxonomy stability across domains, clustering granularities, and LLM configurations, and to further study the robustness of iterative synthesis under different prompting and ordering conditions.
7 Limitations
@@
While our framework reduces manual annotation effort and generalizes across domains, several limitations remain. Firstly, we purposefully choose a smaller LLM to show that our framework is not limited to computational resources or budget constraints. However, the performance of our framework is likely to improve with more powerful LLMs.
Secondly, while we demonstrate the effectiveness of our framework on a political advertisement corpus, it may not generalize to all domains. In particular, domains with highly overlapping or abstract topic boundaries may require additional constraints or hierarchical structure.
@@ A.1 Topic Synthesis
Can the ads can be summarized by one of the previous topics: Answer with "yes" or "no".
@@
In three words or less, describe the issue that ads discuss, summarizing the topic. Examples: "Abortion", "Climate Change", etc.
Output constriained to ["yes", "no"].
@@
A better topic for these ads is: "
The LLM would finish the above, with the quote (") as a stop token.
A.2 Annotation
@@
(a) Females: AL vs WA
(b) All Age groups: FL
(c) Males: MT vs VA
(d) Moral Framing across States
Appendix B Substantive Insights: Topic Salience and Moral Framing
We analyze the distribution of moral framings and political issues across different demographics to uncover political microtargeting strategies. We used the positive point-wise mutual information (PPMI) to detect topics that appear disproportionately in ads shown to specific demographics, potentially indicating targeted messaging. We visualize the results of the PPMI analysis in Figure 4. The heatmaps show the distribution of moral framings and political issues across different demographics. Some interesting observations include the differences in ads targeting different states, or the shift in topics among different age groups in Florida (Figure 4(b)). For example, young people in Florida are more likely to be targeted with ads related to affordable housing, while older people are more likely to be targeted with ads related to abortion. This shows generational targeting strategies, where economic concerns are emphasized for younger groups, while older groups are mobilized on moralized issues like abortion.
Appendix C Substantive Insights: Ad-Level Analysis
Using the synthesized topics, we annotate individual ads to enable ad-level analysis. Cluster-level annotations reveal latent issue and moral structure, while ad-level annotations enable analysis of spending, funder concentration, and targeting strategies.
Figure 5: Top five funding entities by total advertising spend for three major issue areas (Economy, Abortion, and Crime/Justice). The figure shows substantial concentration of spending within each topic, with one or two organizations accounting for a large share of total investment, particularly for abortion and economic messaging, while crime/justice exhibits a comparatively more distributed funding profile across multiple PACs.
C.1 Top Funders
We examine the concentration of advertising spend among funding entities across major issue domains. Fig. 5 presents the top $5$ funding entities by total spend for three central issues in the 2024 election—economy, abortion, and crime/justice [^4] https://www.pewresearch.org/politics/2024/09/09/issues-and-the-2024-election/.
For the economy, spending is highly concentrated among a small number of well-funded national organizations. One entity dominates overall investment, with a sharp drop-off among the remaining top funders. Other contributors—including large advocacy and leadership PACs—account for substantially smaller but still nontrivial shares. This pattern suggests that economic messaging is driven primarily by a limited set of organizations with the capacity to sustain large-scale, sustained advertising campaigns.
The abortion domain exhibits the strongest concentration of spending across the three topics. A single organization accounts for a disproportionate share of total expenditure, while the remaining top funders contribute at comparatively similar but much lower levels. The presence of multiple organizations aligned with opposing positions indicates intense contestation, but the highly skewed spending distribution highlights the outsized role of a small number of actors in shaping abortion-related advertising at scale. This reinforces abortion’s status as a high-salience wedge issue in paid political communication. For example: Florida Voters Against Extremism [^5] https://dos.elections.myflorida.com/committees/ComDetail.asp?account=84315 is the largest spender, anchoring anti-abortion rights mobilization, particularly around state-level ballot initiatives. In contrast, progressive-aligned group such as Montanans Securing Reproductive Rights [^6] https://www.influencewatch.org/organization/montanans-securing-reproductive-rights/ invests heavily in support of abortion rights.
In contrast, the crime and justice topic displays a more distributed funding profile. While two PACs emerge as the largest spenders, the remaining top entities contribute comparable levels of investment, resulting in a less sharply skewed distribution. Several of these organizations are tied to specific states or jurisdictions, indicating a mix of national framing and localized campaign investment. This pattern suggests that crime and justice advertising is both financially intensive and geographically fragmented, reflecting its relevance to state- and local-level electoral contexts.
Overall, Fig. 5 illustrates substantial variation in funding concentration across issue domains. While economy and especially abortion advertising are dominated by a small number of major organizations, crime/justice messaging involves a broader set of actors with more evenly distributed spending, pointing to distinct strategic and organizational dynamics across policy areas.
Figure 6: Aggregate ad spend by moral foundation shows distinct moral structures across topics: abortion centers on liberty/oppression and sanctity/degradation, crime/justice on loyalty/betrayal and authority/subversion, and the economy on fairness/cheating and liberty/oppression.
C.2 Moral Foundations in Funded Messaging
We link moral framing to financial investment across issue domains. Fig. 6 presents total advertising spend by moral foundation for three high-salience topics: abortion, crime/justice, and economy. This aggregate view reveals clear differences in moral structure across issue domains, independent of individual funding entities.
We next examine how these aggregate moral patterns emerge from individual funding entities. Fig. 7 compares the top $5$ funding entities in two salient topics—economy and abortion disaggregated by dominant moral foundation appeals. For the economy, spending is distributed across multiple moral foundations, with substantial investment in both Fairness/Cheating and Liberty/Oppression frames. Several major funders emphasize fairness-oriented rhetoric (e.g., taxation, economic equity), while others concentrate spending almost exclusively on liberty-based appeals, particularly around regulation, government intervention, and market freedom. Minor contributions to Loyalty/Betrayal framing appear for some organizations but remain secondary. Overall, economic messaging exhibits clear moral plurality, with different actors strategically emphasizing distinct moral narratives rather than converging on a single frame.
In contrast, the abortion domain displays a markedly different moral profile. Spending is dominated by two moral foundations—Liberty/Oppression and Sanctity/Degradation with funders sharply specializing in one or the other. Organizations aligned with reproductive rights primarily invest in liberty-based appeals centered on bodily autonomy and government overreach, whereas anti-abortion organizations concentrate heavily on sanctity-based framing. Contributions invoking Care/Harm or Loyalty/Betrayal are present but comparatively limited. This pattern indicates that abortion advertising is characterized not by moral diversity within actors, but by moral polarization across actors.
Taken together, Fig.  7 shows that moral framing is closely tied to both issue domain and funder strategy: economic advertising supports multiple coexisting moral narratives, while abortion advertising is structured around a small number of sharply opposed moral foundations.
Figure 7: Funder-level moral framing for economy and abortion ads: economic messaging spans multiple moral foundations across funders, while abortion advertising is dominated by sharply polarized Liberty/Oppression and Sanctity/Degradation frames.